OpenStudy (anonymous):

what is the diameter of a sphere with volume equal to 14,137.167 cm^3?

OpenStudy (anonymous):

67.380 rounded to the nearest hundredth?

OpenStudy (anonymous):

how did u get that? 14,137.167=4/3pir^3 42411.501=4pir^3 10602.875=pir^3 not sure what to do next